OpenStudy (anonymous):

(2X^2+11X+8) DIVIDED BY (X+4)

OpenStudy (anonymous):

\[(x+4)\times\text{something} = x^2+11x+8\] The something has to have an \(x\) in it to get the \(x^2\) What you can try doing is multiplying out \((x+4)(Ax+B)+C\) and then compare coefficients with \(x^2+11x+8\) For example the coefficient of \(x^2\) is 1, therefore A = 1

OpenStudy (anonymous):

how do I find quotient and remainder?

OpenStudy (anonymous):

The \(Ax+B\) is the quotient, the \(C\) is the remainder

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

I really only need the answer. this is the final question on my test and I don't know how to do it at all

OpenStudy (anonymous):

Ok, well, the way I like to think of these problems (when you're dividing polynomials) is not to think of it as a division, but to think: "What do I need to multiply the thing on the bottom to get the thing on the top?" The thing on the bottom has \(x^1 = x\) as the highest power of \(x\), but the thing on the top has \(x^2\) as the highest power of \(x\). This means that the answer needs and \(x\) and the most general solution will be \[Ax+B+\frac{C}{(x+4)}\] meaning \[(x+4)(Ax+B)+C = x^2+11x+8\] All you need to do is multiply out the left hand side and compare it to the right hand side. So first multiply this out: \[(x+4)(Ax+B)+C = Ax^2+(4A+B)x+4B+C\] Now let's compare coefficients The coefficient of \(x^2\) is \(A = 1\) The coefficient of \(x\) is \(4A+B = 11\) The coefficient of \(1\) is \(4B+C = 8\) Now you can solve for \(A\), \(B\) and \(C\)